OpenStudy (anonymous):

Can someone help me graph the rational function f(x) = quantity x squared plus x minus two divided by quantity x squared minus three x minus four

OpenStudy (anonymous):

factor top and bottom first

OpenStudy (anonymous):

f (x) = x^2 + x - 2/x^2 - 3x - 4 Is factoring like simplifying?

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" you cannot reduce this fraction, but both the top and bottom factor for example \[x^2+x-2=(x+2)(x-1)\]

OpenStudy (anonymous):

that will give you the x intercepts factor the denominator to find the vertical asymptotes

OpenStudy (anonymous):

Would the bottom be x^2 - 3x - 4 = (x + 2)(x - 2)?

OpenStudy (anonymous):

no because \((x+2)(x-2)=x^2-4\)

OpenStudy (anonymous):

I don't understand